# Joint Modelling of Astrophysical Systematics for Cosmology with LSST

### v2

@@ Abstract
We present a novel framework for jointly modelling the weak lensing source galaxy redshift distribution and the intrinsic alignment of galaxies via a shared luminosity function. Considering this framework within the context of a Rubin Observatory’s Legacy Survey of Space and Time (LSST) Year 1 and Year 10 cosmic shear analysis, we first demonstrate the substantial impact of the luminosity function on both source galaxy redshift distributions and the intrinsic alignment contamination. We establish how the individual parameters of a Schechter luminosity function model impact the redshift distribution of a magnitude-limited sample, and we demonstrate the effect of marginalising over the luminosity function parameters as incorporated in the intrinsic alignment modelling of a standard cosmic shear analysis set-up. We forecast the impact of our joint modelling framework on cosmological parameter constraints. Our preliminary results are promising, indicating that this framework can yield cosmological constraints consistent with those expected from standard analyses, enhanced by the flexibility of not fixing luminosity function parameters. We plan to further validate these findings with comprehensive Markov chain Monte Carlo (MCMC) simulations to robustly quantify bias avoidance and underscore the framework’s efficacy. Taking advantage of our forecasting results and in particular the parameter degeneracies which emerge therein, we identify the specific impact of the shape of the luminosity function of source galaxies on the cosmic shear data vector. We also discuss the potential of this method in providing a natural way to model generic selection functions in redshift distribution estimation, as well as its possibilities for extension to a 3$\times$2pt analysis, particularly with respect to incorporating galaxy bias in this luminosity-function-based framework. Although we consider the context of LSST cosmic shear in this work, the proposed joint modelling framework is generically applicable to weak lensing surveys.
@@
gravitational lensing: weak – cosmology: cosmological parameters – Fisher forecast
pubyear: 2024pagerange: Joint Modelling of Astrophysical Systematics for Cosmology with LSST A
@@ 1 Introduction
Weak lensing (WL) is a powerful technique used in cosmology to probe the properties of the universe with high precision. Several Stage III surveys, including the Kilo Degree Survey (Kuijken et al., 2019), the Dark Energy Survey (DES Collaboration et al., 2021), and Hyper Suprime-Cam (Aihara et al., 2022), have already demonstrated its effectiveness in this regard, constraints on specific cosmological parameters comparable to those from the Cosmic Microwave Background (CMB). However, upcoming Stage IV surveys such as the Rubin Observatory’s Legacy Survey of Space and Time (LSST) (Ivezić et al., 2019a), Euclid (Scaramella et al., 2022; Laureijs et al., 2011a), and the Roman Space Telescope (Eifler et al., 2021) promise to further improve the precision of weak lensing measurements. To harness the full potential of these future surveys, it is crucial to develop accurate models for their observables from which we will extract cosmological parameter constraints. An integral component of this task is the modelling of non-cosmological effects, as any inaccuracies in modeling could lead to biases in the cosmological parameter estimation.
@@
Among the significant challenges in modeling non-cosmological effects on weak lensing are astrophysical systematic effects, including the impact of baryonic physics, the galaxy-dark matter bias (for cross-correlations between lensing and galaxy clustering), and, as will be a focus of this work, the intrinsic alignment of galaxies. Intrinsic alignment (IA) refers to the scenario where galaxies exhibit correlated alignment not due to the gravitational lensing of their light but as a result of local effects, predominantly tidal physics as well as galaxy evolution history and environment (Joachimi et al., 2015; Kirk et al., 2015; Troxel & Ishak, 2015). If ignored or incorrectly modelled, the IA impact to the weak lensing cosmological signal has the potential to severely bias cosmological parameter inference (Krause et al., 2015; Secco et al., 2022). Extensive efforts to understand and model intrinsic alignment have thus become a priority, with direct measurement using spectroscopic samples (Mandelbaum et al., 2011; Joachimi et al., 2011; Singh et al., 2015; Johnston et al., 2019), the analysis of alignments in hydrodynamical simulations (Chisari et al., 2015; Tenneti et al., 2016; Samuroff et al., 2021) and theoretical modelling (Hirata & Seljak, 2004b; Bridle & King, 2007; Blazek et al., 2015; Vlah et al., 2020; Maion et al., 2023; Bakx et al., 2023; Chen & Kokron, 2024), all playing a part.
@@
Besides intrinsic alignment modelling, large area photometric survey programs for weak lensing cosmology need to model the redshift distribution of samples of galaxies to unprecedented precision (Mandelbaum et al., 2018). Of particular importance is the accurate and precise quantification of the photometric redshift (PZ) distribution uncertainty, which will require estimation and modelling methodology going beyond that which has been employed in Stage III analyses (see Rau et al. (2023); Bilicki et al. (2021); Myles et al. (2021) for examples of Stage III photometric redshift methodologies and Zhang et al. (2023); Moskowitz et al. (2023); Rau et al. (2020) for examples of new techniques targeted at Stage IV).
@@
The standard technique for mitigating the impact of both IA and PZ uncertainty involves including a model for these effects in the theoretical calculation for weak lensing observables, as part of the parameter inference framework, and marginalising over the parameters of these models when quoting cosmological results (see, e.g. Abbott et al. (2022); Dalal et al. (2023); Giblin et al. (2021)). Currently, these state-of-the-art approaches treat IA and PZ modelling as independent of each other. However, in recent years it has become increasingly recognized that the modelling of intrinsic alignment and photometric redshift distributions within cosmic shear analyses are not fully isolated from each other, with suggestions of trade-offs and interplay between the parameters of the two effects (Wright et al., 2020; Stölzner et al., 2021; Fischbacher et al., 2023). Leonard et al. (2024) recently explored this in the case of a 3$\times$2 point analysis for an LSST-like survey, finding that although biases to cosmological parameters due to model mis-specification of PZ could often be mitigated by marginalisation over a sufficiently flexible IA model, this was not always the case and in general led to significant biases in the IA and PZ parameters themselves. The assumption that IA and PZ uncertainty are independent thus may leave potentially valuable information on the proverbial systematics table.
@@
In this work, we propose a method to link the modelling of IA and PZ uncertainty for cosmic shear, via their mutual dependence on the luminosity of galaxies in the weak lensing source sample. We know that the ensemble redshift distribution in a magnitude-limited sample can be modelled as an integral over a selection function and a luminosity function (see e.g. van Daalen & White 2018). Equally, astrophysical systematic effects including IA, as well as galaxy bias have been shown to be galaxy-luminosity dependent (Fortuna et al., 2020). As shown in van Daalen & White (2018) for the case of galaxy clustering and luminosity-dependent galaxy bias, taking advantage of the link between luminosity-dependent astrophysical systematic effects and luminosity-based modelling of redshift distributions in magnitude limited surveys can provide a natural mechanism for extracting valuable, degeneracy-breaking information.
@@
In this work, we propose a novel modeling framework that establishes a physical connection between the intrinsic alignments and the redshift distribution of galaxies in a magnitude-limited survey. By adopting a shared luminosity function-based approach, we provide a model effective for addressing ensemble redshift distribution uncertainties and selection functions, and we devise a method for self-consistently constraining luminosity function parameters, which are typically fixed based on external datasets. This method facilitates the accurate modeling of selection functions crucial for estimating sample redshift distributions, for instance, in calibration with spatially overlapping spectroscopic data. Extending the work of Joachimi et al. 2011 and Krause et al. 2015, our approach uniquely applies the same luminosity function to simultaneously determine the intrinsic alignment amplitude and the redshift distribution for a magnitude-limited sample. This synergy enhances the physical consistency of our analysis, offering a more integrated and comprehensive understanding of survey observations.
@@
While the focus of this work is primarily on cosmic shear, which quantifies the two-point correlations in the distortion of galaxy shapes due to weak gravitational lensing, the proposed modeling framework has the potential to be extended to 3$\times$2pt analysis and thus incorporate also the luminosity dependence of galaxy bias. The latter involves combining cosmic shear with galaxy-galaxy lensing and galaxy clustering measurements to extract even more valuable cosmological information from the data.
@@
This work is structured as follows: Section 2 (Theory and Methods) provides the theoretical description of key analysis ingredients: the luminosity function, redshift distribution, intrinsic alignments, and the cosmic shear power spectrum. Section 3 describes our forecasting setup. Section 4 presents the results of an exploration of the potential of this method using Fisher forecasting. We conclude in Section 5.
@@ 2.1 Luminosity function of galaxies
Here, $\phi^{*}$ is a normalization factor, $\alpha$ is a power law index, commonly called a faint-end slope, while $L^{*}$ represents a pivot luminosity — a turning point at which the function starts to exponentially decrease from a power law. Since we are interested in the redshift $z$ scaling of the LF, the Schechter function can be expressed as:
$\phi(L,z)=\phi^{*}(z)\left(\frac{L}{L^{*}(z)}\right)^{\alpha}\exp\left(-\frac{L}{L^{*}(z)}\right)\frac{\operatorname{d}\!{L}}{L^{*}(z)}.$ (2)
@@
$\Phi(M)\operatorname{d}\!{M}=0.4\>\phi^{*}\ln(10)\left(10^{0.4(M-M^{*})}\right)^{\alpha+1}\exp\left(-10^{0.4(M-M^{*})}\right)\operatorname{d}\!{M},$ (3)
with $M^{*}$ a characterstic magnitude (corresponding to the pivot luminosity). We have adopted the parametrisation of the LF parameters from Lin et al. (1999) that allows for their redshift evolution:
$\begin{split}\phi^{*}(z)&=\phi_{0}^{*}10^{0.4Pz},\\ M^{*}(z)&=M^{*}_{0}(z_{0})-Q(z-z_{0}),\\ \alpha(z)&=\alpha(z_{0}).\end{split}$ (4)
In this scheme, $Q$ represents a luminosity evolution parameter, describing a linear redshift evolution of characteristic absolute magnitude $M^{*}$. The faint-end slope parameter $\alpha$ is taken to be fixed in redshift in order for the LF shape to be constant. Lastly, $z_{0}$ is the fiducial redshift up to which the magnitudes are color corrected. The other important consequence of this choice is that in this case, the normalization factor $\phi^{*}$ represents the total galaxy number density $\rho=\int\phi(M)\operatorname{d}\!{M}$. This way, the scaling of the total number density of galaxies is regulated by the density evolution parameter $P$, as described in Eq. 4. Furthermore, the redshift evolution of the luminosity density $\rho_{L}=\int L\phi(M)\operatorname{d}\!{M}$ is then conveniently described by evolution parameters $P$ and $Q$:
$\rho_{L}=\rho_{L}(0)10^{0.4(P+Q)z}$ (5)
In constructing the fiducial values of the LF parameters used in this work, we make use as a starting point of parameters from the GAMA survey (Loveday et al., 2012). The GAMA survey based their LF parameter estimation on the maximum-likelihood approach of Lin et al. (1999) for the CNOC2 galaxy redshift survey, but their results differ from those of CNOC2 because of differences in the two samples. We have opted for the GAMA results because of deeper redshift coverage, the completeness in terms of passband filters, and clear distinction between the galaxy types. In this regard, the starting point luminosity-function parameters of our analysis will be $r$-band LF parameters for red and all (red+blue) galaxies (Table 1) with the pivot redshift $z_{0}=0.1$. Details of how we adjust these parameters to find appropriate fiducial luminosity function parameters for our setup are discussed below.
@@ 2.2 Redshift distribution
Arguably, the most fundamental quantity that is necessary for any cosmological and/or astrophysical analysis is an accurate and precise measurement of distances to objects. Cosmological surveys in general employ an approach in which the proxy for the distance (the redshift) is inferred on imaging or spectroscopic data using spectral energy distribution (SED) models. These two approaches are conceptually different methods to study the features of the SED. We can observe large parts of the sky using photometric observations in broad optical filters. However the long exposure times required to measure spectra accurately for faint objects can be prohibitively long (Newman et al., 2015). As a result even with modern multi-object spectrographs, we can only expect to obtain spectra for a few percent of the sources which will be imaged by current and next-generation deep and wide photometric surveys (Salvato et al., 2019).
Since the filters in broad band optical surveys like Rubin are broad and sparse, a lot of the SED information is simply not captured. As a result information on redshift, or distance, is lost by utilizing these lower resolution photometric measurements, as compared with a spectroscopic measurement (up to 100 times, Salvato et al. 2019).
Redshift estimates that have been obtained using photometric data, referred to as photometric redshifts, are commonly employed in cosmological surveys which rely on large data samples, particularly for weak lensing sources.
A photometric redshift (PZ hereafter) estimation method is essentially a mapping between per-band fluxes (hence colors) and the galaxy redshifts. There are several procedures to estimate photometric redshifts, such as template fitting or data driven methods where machine learning (ML) training algorithms are utilized (for a comprehensive review, consult Salvato et al. 2019. Accuracy and precision is demanded by Stage IV surveys, as mentioned above, and understanding the photometric redshift uncertainties are of key importance for achieving these demands. Uncertainties from photometric redshift estimates propagate into the analysis and ultimately alter the cosmological parameter constraints and our understanding of the underlying cosmological model.
@@
Another approach for achieving a redshift distribution estimate was developed by Sheth & Rossi 2010 and further extended by van Daalen & White 2018. This method provides a mechanism for deriving a redshift distribution, denoted as $\frac{\operatorname{d}\!{N}}{\operatorname{d}\!{z}}$, through the process of convolving [^1] We use the term ’convolution’ in a broad sense and not in its strict mathematical definition. It is meant to convey the idea of blending or merging, in this context as summarised in the subsequent equation. a luminosity function $\Phi(M,z)$ with a survey volume $\Delta V(z)$, thereby offering a comprehensive model for understanding galaxy distribution across redshifts.
An illustrative example can be seen in Fig. 1. Fundamentally, the number of galaxies within a given luminosity interval diminishes with increasing redshift, as outlined in Eq. 2 — this is denoted by the solid yellow line in the figure. Conversely, the survey volume, depicted as a teal solid line in Fig. 1, grows with redshift. The result of convolving these two functions yields the redshift distribution of galaxies, represented by the green solid line in the figure.
The enhancements introduced by van Daalen & White (2018) are particularly advantageous for practical applications. This is because the integration now encompasses both redshift $z_{i}$, and magnitude $m_{\lambda}$ bins. The central expression is:
$\langle N_{i,\lambda}\rangle^{\mathrm{Poiss}}=\int^{z_{i,\mathrm{max}}}_{z_{i,\mathrm{min}}}\int^{m_{2}}_{m_{1}}\operatorname{d}\!{m}\operatorname{d}\!{z}\frac{\operatorname{d}\!{\Phi}_{i}(M)}{\operatorname{d}\!{z}}\frac{\operatorname{d}\!{V}}{\operatorname{d}\!{z}}$ (6)
@@
with $\langle N_{i,\lambda}\rangle^{\mathrm{Poiss}}$ the Poisson mean number of galaxies in redshift bin $z_{i}$ and apparent magnitude bin $m_{\lambda}$.
Figure 1: Illustration of the convolution method used in this work. The scaling of the number of galaxies with luminosity as redshift increases, denoted by the solid yellow line (based on Eq. 2). The survey volume, shown as a teal solid line, increases with redshift. The convolution of these two factors produces the galaxy redshift distribution, represented by the green solid line.
The survey volume (in redshift bin $i$) is defined through the comoving distance $D_{\mathrm{C}}$ and the area of the sky covered by the survey $A$ as:
$V_{i}$ $=\int_{A}\int_{D_{i}^{\mathrm{min}}}^{D_{i}^{\mathrm{max}}}\operatorname{d}\!{D}_{\mathrm{C}}(z)\operatorname{d}\!{A}D_{\mathrm{C}}^{2}$
$=4\pi f(A)\int_{z_{i}^{\mathrm{min}}}^{z_{i}^{\mathrm{max}}}\operatorname{d}\!{z}\frac{D_{\mathrm{C}}^{2}(z)c}{H_{0}\sqrt{\Omega_{m,0}(1+z^{\prime})^{3}+\Omega_{\Lambda,0}}}$ (7)
@@
with $f(A)$ being the fraction of the sky, $c$ the speed of light in vacuum, $H_{0}$ the Hubble parameter, and $\Omega_{m,0}$ and $\Omega_{\Lambda,0}$ matter and dark energy density fractions (evaluated at present time), respectively. Flat $\Lambda$CDM framework is assumed.
@@
Crucially, the method described in this subsection enables the direct modelling of the redshift distribution of a magnitude-limited galaxy survey via two physically meaningful ingredients: the luminosity function and the survey volume. Making use of this modelling allows us to connect the modelling of the redshift distribution to other luminosity-dependent effects like intrinsic alignment. It also naturally allows for an extension to incorporate directly within the modelling other survey selection functions beyond a volume selection.

### v3

@@ Abstract
We present a novel framework for jointly modelling the weak lensing source galaxy redshift distribution and the intrinsic alignment (IA) of galaxies through a shared luminosity function (LF). In the context of a Rubin Observatory’s Legacy Survey of Space and Time (LSST) Year 1 and Year 10 cosmic shear analysis, we show that our novel approach produces cosmological parameter constraints which are comparable to standard methods, while offering more physical insight into IA and selection effects. We clarify the relationship between individual parameters of a Schechter luminosity function and the redshift distribution of a magnitude-limited sample, showing the consequences of marginalizing over these parameters when modelling intrinsic alignments in standard cosmic shear analyses. We explore the impact of the shape of the luminosity function on the cosmic shear data vector, and we outline the potential of this method to naturally model selection functions in redshift distribution estimation. Although this work focuses on LSST cosmic shear, the proposed joint modelling framework is broadly applicable to weak lensing surveys.
@@
gravitational lensing: weak – cosmology: cosmological parameters
pubyear: 2024pagerange: Joint Modelling of Astrophysical Systematics for Cosmology with LSST Cosmic Shear A
@@ 1 Introduction
Weak lensing (WL) is a powerful technique used in cosmology to probe the properties of the universe with high precision. Several Stage III surveys, including the Kilo Degree Survey (Kuijken et al., 2019), the Dark Energy Survey (DES Collaboration et al., 2021), and Hyper Suprime-Cam (Aihara et al., 2022), have already demonstrated its effectiveness in this regard, with constraints on specific cosmological parameters comparable to those from the Cosmic Microwave Background (CMB). However, upcoming Stage IV surveys such as the Rubin Observatory’s Legacy Survey of Space and Time (LSST) (Ivezić et al., 2019a), Euclid (Scaramella et al., 2022; Laureijs et al., 2011a), and the Roman Space Telescope (Eifler et al., 2021) promise to further improve the precision of weak lensing measurements. To harness the full potential of these future surveys, it is crucial to develop accurate models for their observables from which we will extract cosmological parameter constraints. An integral component of this task is the modelling of non-cosmological effects, as any inaccuracies in modelling could lead to biases in the cosmological parameter estimation.
@@
Among the significant challenges in modelling non-cosmological effects on weak lensing are astrophysical systematic effects, including the impact of baryonic physics, the galaxy-dark matter bias (for cross-correlations between lensing and galaxy clustering), and, as will be a focus of this work, the intrinsic alignment of galaxies. Intrinsic alignment (IA) refers to the scenario where galaxies exhibit correlated alignment not due to the gravitational lensing of their light but as a result of local effects, predominantly tidal physics as well as galaxy evolution history and environment (Joachimi et al., 2015; Kirk et al., 2015; Troxel & Ishak, 2015). If ignored or incorrectly modelled, the IA impact to the weak lensing cosmological signal has the potential to severely bias cosmological parameter inference (Krause et al., 2016; Secco et al., 2022). Extensive efforts to understand and model intrinsic alignment have thus become a priority, with direct measurement using spectroscopic samples (Mandelbaum et al., 2011; Joachimi et al., 2011; Singh et al., 2015; Johnston et al., 2019), the analysis of alignments in hydrodynamical simulations (Chisari et al., 2015; Tenneti et al., 2016; Samuroff et al., 2021) and theoretical modelling (Hirata & Seljak, 2004; Bridle & King, 2007; Blazek et al., 2015; Vlah et al., 2020; Maion et al., 2023; Bakx et al., 2023; Chen & Kokron, 2024) all playing a part.
@@
Besides intrinsic alignment modelling, large area photometric survey programs for weak lensing cosmology need to model the redshift distribution of samples of galaxies to unprecedented precision (Mandelbaum et al., 2018). Of particular importance is the accurate and precise quantification of the photometric redshift (PZ) distribution uncertainty, which will require estimation and modelling methodology going beyond that which has been employed in Stage III analyses (see Rau et al. 2023; Bilicki et al. 2021; Myles et al. 2021 for examples of Stage III photometric redshift methodologies and Zhang et al. 2023; Moskowitz et al. 2023; Rau et al. 2020 for examples of new techniques targeted at Stage IV).
@@
The standard technique for mitigating the impact of both IA and PZ uncertainty involves including a model for these effects in the theoretical calculation for weak lensing observables, as part of the parameter inference framework, and marginalizing over the parameters of these models when quoting cosmological results (see, e.g. Abbott et al. 2022; Dalal et al. 2023; Giblin et al. 2021). Currently, these state-of-the-art approaches treat IA and PZ modelling as independent of each other. However, in recent years it has become increasingly recognized that the modelling of intrinsic alignment and photometric redshift distributions within cosmic shear analyses are not fully isolated from each other, with suggestions of trade-offs and interplay between the parameters of the two effects (Wright et al., 2020; Stölzner et al., 2021; Fischbacher et al., 2023). Leonard et al. (2024) recently explored this in the case of a 3$\times$2pt analysis for an LSST-like survey, finding that although biases to cosmological parameters due to model mis-specification of PZ could often be mitigated by marginalization over a sufficiently flexible IA model, this was not always the case and in general led to significant biases in the IA and PZ parameters themselves. The assumption that IA and PZ uncertainty are independent thus may leave potentially valuable information on the proverbial systematics table.
@@
In this work, we propose a method to link the modelling of IA and PZ uncertainty for cosmic shear, via their mutual dependence on the luminosity of galaxies in the weak lensing source sample. We know that the ensemble redshift distribution in a magnitude-limited sample can be modelled as an integral over a selection function and a luminosity function (see e.g. van Daalen & White 2018). Equally, astrophysical systematic effects including IA (Blazek et al., 2011; Chisari et al., 2015; Fortuna et al., 2020; Samuroff et al., 2023), as well as galaxy bias have been shown to be galaxy-luminosity dependent. As shown in van Daalen & White (2018) for the case of galaxy clustering and luminosity-dependent galaxy bias, taking advantage of the link between luminosity-dependent astrophysical systematic effects and luminosity-based modelling of redshift distributions in magnitude-limited surveys can provide a natural mechanism for extracting valuable, degeneracy-breaking information.
@@
In this work, we propose a novel modelling framework that exploits the physical connection between the intrinsic alignments and the redshift distribution of galaxies in a magnitude-limited survey. Extending the work of Joachimi et al. (2011), Krause et al. (2016), and van Daalen & White (2018), our approach uses a shared luminosity function to simultaneously model the intrinsic alignment amplitude and the redshift distribution for a magnitude-limited sample. By adopting a shared luminosity function, we provide a method for self-consistently constraining luminosity function parameters (which are typically fixed based on external datasets if included at all), while also introducing a framework for modelling generic selection functions (which often arise when estimating sample redshift distributions, e.g. in calibration with spatially overlapping spectroscopic data). Although in this initial proof-of-concept work we incorporate photometric redshift information only via the impact of the redshift distribution on the weak lensing two-point datavector, our method lays the groundwork for a more general unified framework to incorporate a single luminosity function within both modelling of astrophysical systematics and estimation of redshift distributions via spectral fitting and cross-correlation methods.
@@
While the focus of this work is primarily on cosmic shear, which quantifies the two-point correlations in the distortion of galaxy shapes due to weak gravitational lensing, the proposed modelling framework has the potential to be extended to 3$\times$2pt analysis and thus incorporate also the luminosity dependence of galaxy bias. The latter involves combining cosmic shear with galaxy-galaxy lensing and galaxy clustering measurements to extract even more valuable cosmological information from the data.
@@
This work is structured as follows: Section 2 provides the theoretical description of key analysis ingredients: the luminosity function, redshift distribution, intrinsic alignments, and the cosmic shear power spectrum. Section 3 describes our forecasting setup. Section 4 presents the results of an exploration of the potential of this method using Fisher forecasting. We conclude in Section 5.
@@ 2.1 Luminosity function of galaxies
Here, $\phi^{*}$ is a normalization factor, $\alpha$ is a power law index, commonly called a faint-end slope, while $L^{*}$ represents a pivot luminosity — a turning point at which the function starts to exponentially decrease from a power law. Since we are interested in the redshift, $z$, scaling of the LF, the Schechter function can be expressed as:
$\phi(L,z)dL=\phi^{*}(z)\left(\frac{L}{L^{*}(z)}\right)^{\alpha}\exp\left(-\frac{L}{L^{*}(z)}\right)\frac{\operatorname{d}\!{L}}{L^{*}(z)}.$ (2)
@@
$\begin{split}\Phi(M,z)\operatorname{d}\!{M}&=0.4\>\phi^{*}(z)\ln(10)\left(10^{0.4(M-M^{*}(z))}\right)^{\alpha+1}\\ &\times\exp\left(-10^{0.4(M-M^{*}(z))}\right)\operatorname{d}\!{M},\end{split}$ (3)
with $M^{*}(z)$ a characteristic magnitude (corresponding to the pivot luminosity). We have adopted the classic but still widely used parametrization of the LF parameters from Lin et al. (1999) (see e.g. Loveday et al. 2012, Farrow et al. 2015, and Karademir et al. 2023 for examples of usage). Under this choice, the redshift evolution of the Schechter parameters is given by:
$\begin{split}\phi^{*}(z)&=\phi_{0}^{*}10^{0.4Pz},\\ M^{*}(z)&=M^{*}_{0}|_{z_{0}}-(z-z_{0})Q,\\ \alpha(z)&=\alpha|_{z_{0}}.\end{split}$ (4)
In this scheme, $Q$ is a scalar that represents a luminosity evolution parameter, describing a linear redshift evolution of characteristic absolute magnitude $M^{*}$. The faint-end slope parameter $\alpha$ is taken to be fixed in redshift in order for the LF shape to be constant. Lastly, $z_{0}$ is a fixed pivot redshift. An important consequence of this choice is that in this case, the normalization factor $\phi^{*}(z)$ is closely related to the total galaxy number density $\rho=\int\phi(M)\operatorname{d}\!{M}$. This way, the scaling of the total number density of galaxies is regulated by the density evolution parameter $P$, as described in Eq. 4. Furthermore, the redshift evolution of the luminosity density $\rho_{L}=\int L\phi(M)\operatorname{d}\!{M}$ is then conveniently described by evolution parameters $P$ and $Q$:
$\rho_{L}=\rho_{L}(0)10^{0.4(P+Q)z}.$ (5)
In constructing the fiducial values of the LF parameters used in this work, we make use as a starting point the parameters from the GAMA survey (Loveday et al., 2012). The GAMA survey based their LF parameter estimation on the maximum-likelihood approach of Lin et al. (1999) for the CNOC2 galaxy redshift survey, but their results differ from those of CNOC2 because of differences in the two samples. We have opted for the GAMA results because of deeper redshift coverage, the completeness in terms of passband filters, and clear distinction between the galaxy types. In this regard, the starting point luminosity-function parameters of our analysis will be $r$-band LF parameters for red and all (red+blue) galaxies (Table 1) with the pivot redshift $z_{0}=0.1$. Details of how we adjust these parameters to find appropriate fiducial luminosity function parameters for our setup are discussed below.
@@ 2.2 Redshift distribution
Arguably, the most fundamental quantity that is necessary for any cosmological and/or astrophysical analysis is an accurate and precise measurement of distances to objects. Cosmological surveys in general employ an approach in which the proxy for the distance (the redshift) is inferred on imaging or spectroscopic data using spectral energy distribution (SED) models. These two approaches are conceptually different methods to study the features of the SED. We can observe large parts of the sky using photometric observations in broad optical filters. However, the long exposure times required to measure spectra accurately for faint objects can be prohibitively long (Newman et al., 2015). As a result, even with modern multi-object spectrographs, we can only expect to obtain spectra for a few percent of the sources which will be imaged by current and next-generation deep and wide photometric surveys (Salvato et al., 2019). Since the filters in broad-band optical surveys like Rubin are broad and sparse, a lot of the SED information is simply not captured. As a result information on redshift, or distance, is lost by utilizing these lower resolution photometric measurements, as compared with a spectroscopic measurement (up to 100 times, Salvato et al. 2019).
Redshift estimates that have been obtained using photometric data, referred to as photometric redshifts, are commonly employed in cosmological surveys which rely on large data samples, particularly for weak lensing sources. An SED-based photometric redshift (PZ hereafter) estimation method is essentially a mapping between per-band fluxes (hence colors) and the galaxy redshifts. There are several procedures to estimate photometric redshifts, such as template fitting or data-driven methods where machine learning (ML) training algorithms are utilized (for a comprehensive review, consult Salvato et al. 2019). Accuracy and precision is demanded by Stage IV surveys, as mentioned above, and understanding the photometric redshift uncertainties are of key importance for achieving these demands. Uncertainties from photometric redshift estimates propagate into the analysis and ultimately alter the cosmological parameter constraints and our understanding of the underlying cosmological model.
@@
Another approach for achieving a redshift distribution estimate was developed by Sheth & Rossi (2010) and further extended by van Daalen & White (2018). This method provides a mechanism for deriving a redshift distribution, denoted as $\frac{\operatorname{d}\!{N}}{\operatorname{d}\!{z}}$, via the appropriate combination of a luminosity function $\Phi(M,z)$ with a redshift-dependent survey volume element. We now introduce this method in more detail.
The central expression is (van Daalen & White, 2018):
$\langle N_{i,\lambda}\rangle^{\mathrm{Poiss}}=\int^{z_{i,\mathrm{max}}}_{z_{i,\mathrm{min}}}\operatorname{d}\!{z}\int^{m_{\lambda,\mathrm{max}}}_{m_{\lambda,\mathrm{min}}}\operatorname{d}\!{m}\frac{\operatorname{d}\!{\Phi}(M(m,z),z)}{\operatorname{d}\!{z}}\frac{\operatorname{d}\!{V}_{C}(z)}{\operatorname{d}\!{z}}$ (6)
@@
with $\langle N_{i,\lambda}\rangle^{\mathrm{Poiss}}$ the Poisson mean number of galaxies in redshift bin $i$ and apparent magnitude bin $\lambda$. $z_{i,\mathrm{min}}$ and $z_{i,\mathrm{max}}$ are the lower and upper limits of redshift bin $i$ respectively, and similar for $m_{\lambda,\mathrm{min}}$ and $m_{\lambda,\mathrm{max}}$ with respect to apparent magnitude bin $\lambda$.
Figure 1: Illustration of the approach used to model the redshift distribution of source galaxies in this work. The luminosity function as calculated using Eq. 2 with fiducial parameters is shown in dashed yellow (evaluated for this illustrative example at $m=17.4$ for the purpose of visualizing the redshift dependence). The derivative with respect to redshift of the comoving survey volume, $V_{C}$ (as can be found using Eq. 2.2), is shown as a short-dashed teal line. Combining these two quantities as in Eq. 6 produces the galaxy redshift distribution, represented by the green solid line and normalized here such that it peaks at unity.
Eq. 6 is constructed by integrating the number density (as related to the redshift-dependent luminosity function) over the comoving volume element of the survey on the appropriate magnitude and redshift range (for details, see van Daalen & White 2018). The comoving survey volume element, $d\,V_{C}$, enters Eq. 6 via the $\frac{\operatorname{d}\!{V}_{C}(z)}{\operatorname{d}\!{z}}$ term. It is defined through the comoving distance $D_{\mathrm{C}}(z)$ and the area of the sky covered by the survey $A_{\mathrm{sky}}$:
$d\,V_{C}$ $=A_{\mathrm{sky}}\,D_{\mathrm{C}}^{2}(z)dD_{\mathrm{C}}$
$=4\pi f_{\mathrm{sky}}\frac{D_{\mathrm{C}}^{2}(z)c}{H_{0}\sqrt{\Omega_{m,0}(1+z)^{3}+\Omega_{\Lambda,0}}}\operatorname{d}\!{z}$
$=\frac{dV_{C}(z)}{dz}dz$ (7)
@@
with $f_{\mathrm{sky}}$ being the fraction of the sky, $c$ the speed of light in vacuum, $H_{0}$ the Hubble parameter, and $\Omega_{m,0}$ and $\Omega_{\Lambda,0}$ matter and dark energy density fractions (evaluated at present time), respectively. A flat $\Lambda$CDM framework is assumed.
An illustration of the application of Eq. 6 can be seen in Fig. 1. Fundamentally, the number of galaxies within a given luminosity interval diminishes with increasing redshift, as outlined in Eq. 2 — this is shown by the long-dashed yellow line (for a fixed apparent magnitude for illustrative purposes). Conversely, the derivative of the survey volume with respect to $z$, which is depicted as the short-dashed teal solid line in Fig. 1, grows with redshift. The result of these two behaviors yields, via Eq. 6, the redshift distribution of galaxies, represented by the green solid line.
@@
The method described in this subsection enables the direct modelling of the redshift distribution of a magnitude-limited galaxy survey via two physically meaningful ingredients: the luminosity function and the survey volume. Making use of this modelling allows us to connect the modelling of the redshift distribution to other luminosity-dependent effects like intrinsic alignment. It also naturally allows for an extension to incorporate directly within the modelling other survey selection functions beyond a volume selection.